OpenStudy (anonymous):

Use any problem solving strategy to solve the following problem. suppose you are fencing a rectangular dog pen with a 100ft of fencing. The pen is to be 10 feet longer than it is wide. what are the dimensions of the pen?

OpenStudy (anonymous):

let x be the width, then x+10 will be its length the perimeter of the pen is 100 = 2x + 2(x+10) solve for x: 100 = 2x + 2x +20 80 = 4x x = 20 x+10 = 30 so the width = 20 ft and length = 30 ft

OpenStudy (anonymous):

100 ft of fencing means that the total perimeter of the fence is 100ft. Now if you take the width of the fence to be x then the length x + 10

OpenStudy (anonymous):

The formula for a rectangle is P = 2w + 2l so 100 = 2x + 2(x + 10) 100 = 2x + 2x + 20...distribute the 2 100 = 4x + 20....Add like terms (2x + 2x = 4x) 4x = 80....subtract 20 from both sides x = 20...divide both sides by 4. Pen is x + 10 = 20 + 10 = 30ft